OpenStudy (anonymous):

translate to a system of equations and solve Molly is 4 years is younger than heidi in 3 years Heidi will be ticw as old as Molly. how old are they now?

OpenStudy (anonymous):

\[Molly + 4 = Heidi\] \[2*(Molly + 3) = Heidi + 3\] Can you solve it on your own?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

you know from the first statement/equation, that Heidi = Molly + 4 . I you substitute this into the second Equation it looks like this: \[2 * (Molly + 3) = (Molly + 4) + 3\] So now you can solve for Molly and plug it back into either the first or the second equation...

OpenStudy (anonymous):

so it is 6 and 11

OpenStudy (radar):

I don't think so, because it was said that Molly is 4 years younger than Heidi, and 11-6=5 should be 4!!

OpenStudy (anonymous):

then how old are they now??

OpenStudy (radar):

You got Heidi's age correct. Review your work and see if you made an arithmetical error.

OpenStudy (radar):

Actually you did not get heidi's age right. They are both wrong.

OpenStudy (radar):

Do this one over and see what you get. 2 * (Molly + 3)= (Molly +4)+3 What did you get for Molly?

OpenStudy (radar):

Let M= Mollly, so that it will look like this: 2(M+3)=(M+4)+3

OpenStudy (radar):

Solve for M

OpenStudy (radar):

The solution is for their age now.

OpenStudy (anonymous):

2 and 6???

OpenStudy (radar):

Lets see where you are going with this. starting with the eequation 2(M+3)=(M+4)+3 Remove paren' 2M+6=M+4+3 consolidating like terms 2M+6=M+7 do we agree thus far?

OpenStudy (radar):

Next step is to subtract 1M and 6 from both sides of the equal sign. 2M-M+6-6=M-M+7-6 Consolidating similar terms Now you have Molly's age!

OpenStudy (radar):

What did you get when you combined the terms of the equation?

OpenStudy (radar):

Have you reached the point where you discovered your previous error?

OpenStudy (anonymous):

so it is one and five?????

OpenStudy (radar):

Absolutely it checks out.

OpenStudy (radar):

Anything about the solution method that you haves questions or comments ??

OpenStudy (anonymous):

it is 1 and 5??? did i get it right???

OpenStudy (anonymous):

yes, it it - you got the right answer.